# Chart-R1: Chain-of-Thought Supervision and Reinforcement for Advanced Chart Reasoner

### v2

@@ 4 Experiments
4.1 Implementation Details
For data generation, we employ Gemini-2.5-Flash to create both plotting code and QA pairs. In the training stage, our ChartRQA-SFT is used for SFT, while a combination of ChartQA and ChartRQA-RL is used for GRPO. The SFT stage is trained for one epoch with a batch size of 48, and the RL stage is trained for 3 epochs with a batch size of 128. The final Chart-R1 model is obtained by applying this RL process to the initial SFT-trained model, Chart-R1-SFT. For these respective stages, the learning rates are set to 1e-5 and 1e-6. Finally, the training processes for SFT and RL required approximately 3 and 30 hours, respectively, on a system with 24 H800 GPUs.
Table 3: The main results on existing chart understanding and reasoning benchmarks. Our proposed Chart-R1 outperforms all the small-scale VLMs (<20B) on the evaluation benchmarks. Bold denotes the best performances of open-source VLMs.
Model Name ChartQA CharXiv-RQ ChartQAPro ChartRQA (single / multi)
Proprietary
GPT-4o 85.7 47.1 37.67 44.37 / 46.55
Gemini-1.5-Flash 79.0 33.9 42.96 -
Gemini-1.5-Pro 87.2 43.3 - -
Gemini-2.5-Flash - - - 59.12 / 59.17
Claude-3.5-Sonnet 90.8 60.2 43.58 52.79 / 56.05
General-domain Open-source
Phi-3.5-Vision 81.8 32.7 24.73 31.08 / 24.32
DeepSeek-VL2 86.0 - 16.28 23.15 / 20.29
InternVL3-8B 86.6 37.6 - 37.51 / 31.73
InternVL3-38B 89.2 46.4 - 46.09 / 38.36
Qwen2.5-VL-7B 87.3 42.5 36.61 44.59 / 40.57
Chart-domain
ChartLlama 69.66 14.2 - -
TinyChart 83.60 8.3 13.25 6.75 / 6.11
ChartGemma 80.16 12.5 6.84 7.18 / 9.23
ChartReasoner 86.93 - 39.97 -
Chart-R1-7B (Ours) 91.04 46.2 44.04 52.09 / 49.93
@@ 4.2 Experiment Settings
We conduct experiments to evaluate the results obtained from various training settings. Firstly, we assess the training stages and the scope of training data, including: (1) SFT with CoT data, (2) Directly RL versus CoT-RL, and (3) RL w/ and w/o the ChartRQA data.
Benchmarks To comprehensively evaluate the understanding and reasoning capacity of our posed Chart-R1, we choose ChartQA Masry et al. (2022), Chaxiv-RQ (Reasoning Questions) Wang et al. (2024), ChartQAPro Masry et al. (2025) and our proposed ChartRQA as the evaluation benchmarks.
Baselines We compare our proposed Chart-R1 with existing models in three setups: (1) Proprietary models include GPT-4oOpenAI (2024), Gemini-1.5-(Flash, Pro)Team et al. (2023), Gemini-2.5-Flash and Claude-3.5-SonnetAnthropic (2024). (2) General-domain open-source VLMs including Phi 3.5-Vision Abdin et al. (2024), DeepSeek-VL2 Wu et al. (2024), InternVL3(8B, 38B) Zhu et al. (2025) and Qwen2.5-VL(7B) Bai et al. (2025). (3) Chart-domain VLMs including ChartLlama Han et al. (2023), TinyChart Zhang et al. (2024), ChartGemma Masry et al. (2024b) and ChartResoner Jia et al. (2025).
@@ 4.3 Main Results
Table 3 show the performance of Chart-R1 compared with other baseline models. The results show that Chart-R1 achieve the state-of-the-art performance on small-scale (<20B) VLMs, including general- and chart-domain models across all the benchmarks. Especially in ChartQA, Chart-R1 achieves the best performance, even compared with proprietary and large-scale VLMs. In the chart reasoning benchmark, CharXiv-RQ, ChartQAPro and our proposed ChartRQA, Chart-R1 significantly surpass existing chart-domain models. Since the training data of Chart-R1 only contains ChartRQA and ChartQA, these results demonstrate the diversity of our proposed ChartRQA dataset and CoT-RL training strategy.
@@
Figure 3: The training curve during the RL stage that utilizing the ChartQA dataset solely.
Figure 4: Training curves for the RL stage using the ChartQA and ChartRQA datasets. The orange curve represents our proposed two-stage training strategy, while the blue curve corresponds to a baseline RL-only setting.
@@ 4.4 Ablation Study
We first assess the impact of different training settings, with results presented in Table 4. The findings indicate that utilizing our two-stage training strategy, Chart-SFT data for step-by-step SFT and ChartQA and ChartRQA-RL data for RL yields the most balanced performance. Notably, omitting Chart-COT causes a significant performance drop on the ChartRQA benchmark. We attribute this to the nature of ChartRQA as complex charts requiring multi-step thinking before answering. The first Chart-COT stage equips the model with the necessary capability for such step-by-step task decomposition. Also, SFT exclusively on the ChartQA dataset leads to performance degradation across all benchmarks, including ChartQA itself. We reckon that although SFT could improve capacity for in-domain tasks, training on simple and low-diversity datasets disrupts the tuned distribution, harming the ability on both in-domain (ChartQA) and OOD (CharXiv-RQ, ChartRQA) tasks.
@@
Previous research has established that the complexity of the training data is critical for effective RL Guo et al. (2025). Our generated ChartRQA training set meets this requirement, featuring tasks with single- and multi-chart images and questions requiring step-by-step reasoning. Including our ChartRQA dataset during the RL stage is crucial for achieving optimal performance. The structural and logical complexity is important for performance enhancements are observed in our Chart-RFT stage. We find that training exclusively on the ChartQA dataset is insufficient for developing a robust reasoning model. The limited complexity of ChartQA fails to encourage the model to learn diverse, long-path reasoning strategies. This limitation is empirically demonstrated by the training process shown in Figure 3. The accuracy reward rapidly converges to around 0.9 with little subsequent growth, while the response length remains constrained to approximately 100 tokens.
Table 5: Ablation study on the accuracy reward. The results show that utilizing specialized reward functions for different task types achieves superior performance.
Accuracy Reward ChartQA CharXiv-RQ ChartRQA
Edit Distance Soft Matching (single / multi)
89.88 44.0 45.02 / 39.79
90.28 45.2 44.16 / 40.44
Figure 5: The visualization result of a case study that Chart-R1 (w/ and w/o Chart-COT) answer correctly, but Qwen2.5VL-7B fails.
Table 6: Ablation study about the combination of SFT data.
SFT Training Set ChartQA CharXiv-RQ ChartRQA (single / multi)
RQA-SFT 89.88 44.5 48.02 / 48.11
RQA-SFT + QA-Train+ RQA-RL 88.40 41.2 45.98 / 44.60
@@
We further investigate the impact of our two-stage training strategy, which involves the Chart-COT and Chart-RFT stages. By comparing to a baseline without the Chart-COT phase, we analyze the resulting training process, specifically the reward and response length. We find that the first SFT on CoT data has two key benefits. First, it significantly increases the token length generated during the RL phase. Second, it leads to a much effective accuracy reward curve, which rises quickly at the start of training and then converges at a higher final value.
Reward Function For the reward function, we conduct experiments to assess the different settings of the accuracy reward. The results are denoted in Table 5. To better assess the influence caused by different accuracy rewards, we conduct ablation studies that train Qwen2.5VL-7B-Instruct for the RL stage only. The results demonstrate that employing a soft accuracy reward, which combines edit distance for string-based tasks and soft matching for numerical tasks, yields superior performance across the majority of our benchmarks. This finding underscores the importance of adjusting the reward function to the specific type of answers.
SFT Data When training Chart-R1, our SFT dataset consists of 228k samples from our ChartRQA-SFT. We then ablate the SFT data composition by adding two sources, the ChartQA dataset and the 30k ChartRQA-RL that overlaps with the RL data, to assess the impact on performance. We train each setting for 2k steps and 1 epoch for SFT and RL, respectively. The results in Table 6 show that combining ChartQA and ChartRQA-RL, the final performance decreases evidently. Our analysis indicates that using overlapping data for SFT and RL leads to overfitting, where the model memorizes reasoning paths from the SFT stage, resulting in more rigid thinking processes and a significant loss of output diversity. Also, the direct-answer format of the ChartQA data discourages the model from developing the ability to break down problems into a step-by-step thinking process.
@@
Visualization We conclude with qualitative case studies where our Chart-R1 model successfully generates detailed reasoning and correct answers for complex questions in Figures 5 and 6. In these same instances, the baseline Qwen2.5VL-7B model fails, directly demonstrating the superior performance and more advanced reasoning capabilities of our approach. Without training on the Chart-COT stage, Chart-R1 w/o Chart-COT also fails to answer problems in Figure 6, which shows the importance of our proposed two-stage training.
@@
Figure 6: The visualization result of a case study that Chart-R1 (w/ Chart-COT) answers correctly, but Qwen2.5VL-7B and Chart-R1 (w/o Chart-COT) fail.
@@ 5 Conclusion
In this paper, we propose Chart-R1, a chart-domain VLM for complex chart reasoning. To improve the reasoning capacity of Chart-R1, we introduce a programmatic data generation method alongside a novel two-stage training strategy to optimize the data construction and training methodology. Also, we propose ChartRQA, which contains 258k training samples which each constructed in verifiable formats and a benchmark for evaluating complex chart reasoning. The result shows that combining our proposed training strategy, Chart-R1 achieves superior performance compared with other VLMs.
Appendix A More Training Details
In this section, we provide more details about Chart-R1’s training process, including Chart-COT and Chart-RFT.
Chart-COT We use Qwen2.5VL-7B-Instruct as the initial model and perform supervised fine-tuning using LLaMA-Factory Zheng et al. (2024). We train the model on the 228k ChartRQA-SFT dataset for one epoch. During training, we freeze the vision tower and multi-modal projector parameters and tune the LLM. The learning rate is set to 1e-5, with a warm-up ratio of 0.1 and batch size of 48. The training process costs 3 hours on 24 H800 GPUs.
Chart-RFT For the RFT stage, we use the fine-tuned model from the Chart-COT stage. We adopt the MM-EUREKA Meng et al. (2025a) framework based on OpenRLHF for training. The model is trained for 3 episodes using 30k ChartQA and 30k ChartRQA-RL. We set the rollout batch size and the training batch size to 128, with each sample generating 8 rollouts. The temperature for model generation is set to 1, and we exclude KL divergence in the loss calculation. The learning rate is set to 1e-6, with a warm-up ratio of 0.03, while freezing the vision tower during training. Following the default setting for instruction models, the format reward coefficient is set to 0.5. We employ the online filtering strategy with lower and upper bounds of 0.1 and 0.9, respectively. The training process costs 30 hours on 24 H800 GPUs.
Appendix B Benchmark Details
@@
ChartQAPro Masry et al. (2025) introduces a diverse benchmark with various chart types, including infographics and dashboards, and question formats that better reflect real-world challenges. We evaluate models using Chain-of-Thought (CoT) prompting in the original paper and report overall accuracy across five question types.
Appendix C ChartRQA Analysis
We count the quantity and distribution of different chart types across the training and test sets of ChartRQA, as detailed in Table A. The distribution among the various types to be well-balanced. Furthermore, Figures A and  B provide visualization examples of 24 chart types from the ChartRQA dataset, showcasing both single-chart and multi-chart formats, respectively.
Table A: The detailed chart types and corresponding quantities in our proposed ChartRQA train and test set. ChartRQA contains 24 chart types, each of which contains approximate samples.
@@ Appendix C ChartRQA Analysis
Figure A: Single-chart samples of 24 chart types from ChartRQA.
Figure B: Multi-chart samples of 24 chart types from ChartRQA.
Appendix D Prompts
To enhance transparency and reproducibility, we provide the exact prompts used for dataset generation and evaluation.
Figure C illustrates the prompt used for code generation. We utilize real table data as input, select one chart type from the 24 predefined chart types, and sample a code example corresponding to that chart type to generate the plotting code for the given table data.
Figures D and  E display the prompts used to generate reasoning QA pairs for single-chart and multi-chart formats, respectively. We craft an example for each format to aid LLMs in understanding complex chart reasoning tasks and to generate step-by-step reasoning processes and precise answers that conform to the format. The executable plotting code is provided as auxiliary information to LLMs, making the generated QA pairs more reliable.
Figure F shows the prompt used for model evaluation. We employ GPT-4o to assess the match between the ground truth and the model’s predictions, where GPT-4o returns a score of 0 or 1 to indicate the correctness of the model’s prediction. Our evaluation focuses solely on the correctness of the final answer, disregarding the reasoning process.
Figure C: Prompt for code generation.
Figure D: Prompt for reasoning QA pairs generation for single-chart formats.
Figure E: Prompt for reasoning QA pairs generation for multi-chart formats.
Figure F: Prompt for ChartRQA evaluation using GPT-4o.

### v3

@@ 4.1 Experiment Settings
We conduct experiments and ablation studies to evaluate the results obtained from various training settings. See the appendix for training details.
Benchmarks To comprehensively evaluate the understanding and reasoning capacity of Chart-R1, we choose ChartQA Masry et al. (2022), CharXiv-RQ (Reasoning Questions) Wang et al. (2024), ChartQAPro Masry et al. (2025a), ChartMuseum Tang et al. (2025a) and our proposed ChartRQA (single/multi) as the evaluation benchmarks.
Baselines We compare our proposed Chart-R1 with existing models in three setups: (1) Proprietary models include GPT-4o, GPT-4.1 OpenAI (2025a), Gemini-1.5-(Flash, Pro), Gemini-2.5-Flash Kavukcuoglu (2025), Claude-3.5-Sonnet and Claude-3.7-Sonnet Anthropic (2025). (2) General-domain open-source VLMs, including Phi 3.5-Vision Abdin et al. (2024), DeepSeek-VL2 Wu et al. (2024), InternVL3(8B, 38B) Zhu et al. (2025) and Qwen2.5-VL(7B) Bai et al. (2025); and VLM reasoning models, including R1-VL Zhang et al. (2025), VL-Rethinker Wang et al. (2025a) and MMR1 Leng et al. (2025). (3) Chart-domain VLMs including ChartLlama Han et al. (2023), TinyChart Zhang et al. (2024), ChartGemma Masry et al. (2024b), ChartResoner Jia et al. (2025), BigCharts-R1 Masry et al. (2025b) and Bespoke-MiniChart-7B Tang et al. (2025b).
Figure 3: The training curve during the RL stage that utilizes the ChartQA dataset solely.
Figure 4: Training curves for the RL stage using the ChartQA and ChartRQA datasets. The orange curve represents our proposed two-stage training strategy, while the blue curve corresponds to a RL-only baseline.
@@ 4.2 Main Results
Table 3 shows the performance of Chart-R1 compared with other baseline models. The results show that Chart-R1 achieves the state-of-the-art performance on small-scale (<<20B) VLMs, including general- and chart-domain models across various chart understanding and reasoning benchmarks. Especially in ChartQA, Chart-R1 achieves the best performance, even compared with proprietary and large-scale VLMs. In the chart reasoning benchmark, CharXiv-RQ, ChartMuseum and our proposed ChartRQA, Chart-R1 significantly surpass existing chart-domain models. Since the training data of Chart-R1 only contains ChartRQA and ChartQA, these results demonstrate the effectiveness of our proposed ChartRQA dataset and CoT-RL training strategy.
@@ 4.3 Ablation Study
We first assess the impact of different training settings, with results presented in Table 4. The findings indicate that utilizing our two-stage training strategy yields the most balanced performance. Notably, omitting Chart-COT causes a significant performance drop on the ChartRQA benchmark. We attribute this to complex charts requiring multi-step thinking before answering. The first Chart-COT stage equips the model with the necessary capability for such step-by-step task decomposition. Also, SFT on the ChartQA dataset leads to performance degradation across all benchmarks, including ChartQA itself. We reckon that although SFT could improve capacity for in-domain tasks, training on simple and low-diversity datasets disrupts the tuned distribution, harming the ability on both in-domain (ChartQA) and OOD (CharXiv-RQ, ChartRQA) tasks.
@@
Prior research underscores the critical role of training data complexity for effective RL Guo et al. (2025). Our generated ChartRQA training set addresses this by featuring tasks with both single- and multi-chart images, and questions demanding step-by-step reasoning. Including our ChartRQA dataset during the RL stage is crucial for achieving optimal performance. The structural and logical complexity is important for performance enhancements observed in our Chart-RFT stage. Furthermore, RL exclusively on the ChartQA dataset is insufficient for developing a reasoning model. The limited complexity of ChartQA fails to encourage the model to learn diverse, long-path reasoning strategies. This limitation is empirically demonstrated by the training process shown in Figure 3. The accuracy reward rapidly converges to around 0.9 with little subsequent growth, while the response length remains constrained to approximately 100 tokens.
@@
We further investigate the impact of our two-stage training strategy, comparing it to a baseline without the Chart-COT phase. The comparison of RL processes is shown in Figure 4. We find that the first SFT on CoT data has two key benefits. First, it significantly increases the token length generated during the RL phase. Second, it leads to a much effective accuracy reward curve, which rises quickly at the start of training and then converges at a higher final value.
@@
Visualization We present qualitative case studies where our Chart-R1 model successfully generates detailed reasoning and correct answers for complex questions in Figure 5. In these same instances, the baseline Qwen2.5-VL-7B model fails, directly demonstrating the superior performance and more advanced reasoning capabilities of our approach. When Chart-R1 is trained without the Chart-COT stage, it also fails to answer the problems in the right case of Figure 5. Although it can correctly recognize the chart content, it makes errors during the reasoning process, highlighting the importance of our proposed two-stage training.
@@
Figure 5: Visualization results of two case studies: (a) Chart-R1 (with and without Chart-COT) both answer correctly while Qwen2.5-VL-7B fails, and (b) only Chart-R1 with Chart-COT answers correctly while both Qwen2.5-VL-7B and Chart-R1 without Chart-COT fail.
4.4 Error Analysis
Chart-R1 achieves significant improvements in reasoning ability compared to baseline, but there is still room for further improvement. We randomly sample 50 incorrect responses from Chart-R1 on ChartQAPro and analyze the error types. As shown in Figure 6, Chart-R1 is most prone to errors in visual reasoning, multi-chart QA, and unanswerable types. Visual reasoning is more challenging than mathematical reasoning, as the latter mainly involves numerical recognition and calculation, while the former requires the model to identify and summarize complex chart patterns. Multi-chart QA requires the model to integrate information across multiple charts. While ChartRQA was designed to address multi-chart reasoning, the current model still exhibits deficiencies in this aspect. For unanswerable questions, although ChartRQA did not specifically include such samples, Chart-R1 can reject most unanswerable questions through RL, demonstrating good generalizability.
Figure 6: Error type distribution of Chart-R1 on ChartQAPro.
@@ 5 Conclusion
In this paper, we propose Chart-R1, a chart-domain VLM for complex chart reasoning. To improve the reasoning capacity of Chart-R1, we introduce a programmatic data generation method alongside a novel two-stage training strategy to optimize the data construction and training methodology. Also, we propose ChartRQA, which contains 258k training samples, each constructed in verifiable formats and a benchmark for evaluating complex chart reasoning. The result shows that combining our proposed training strategy, Chart-R1 achieves superior performance compared with other VLMs.
Limitations
Our study focuses primarily on statistical charts from academic papers, overlooking practical visualization types such as dashboards and flowcharts. This leads to a gap compared to closed-source models on comprehensive benchmarks such as ChartMuseum. In future research, we plan to expand our training paradigm to incorporate diverse chart types and complex visual reasoning, developing a more versatile chart understanding model.
Ethical Statement
Our research employs publicly available models and datasets with proper citations. This approach leverages the widespread use and non-toxic nature of our datasets and prompts to minimize the risk of generating toxic content.
Appendix A Appendix
A.1 Training Details
Chart-COT We use Qwen2.5-VL-7B-Instruct as the initial model and perform supervised fine-tuning using LLaMA-Factory Zheng et al. (2024). We train the model on the 228k ChartRQA-SFT dataset for one epoch. During training, we freeze the vision tower and multi-modal projector parameters and tune the LLM. The learning rate is set to 1e-5, with a warm-up ratio of 0.1 and batch size of 48. The training process costs 3 hours on 24 H800 GPUs.
Chart-RFT For the RFT stage, we use the fine-tuned model from the Chart-COT stage. We adopt the MM-EUREKA Meng et al. (2025b) framework based on OpenRLHF for training. The model is trained for 3 episodes using 30k ChartQA and 30k ChartRQA-RL. We set the rollout batch size and the training batch size to 128, with each sample generating 8 rollouts. The temperature for model generation is set to 1, and we exclude KL divergence in the loss calculation. The learning rate is set to 1e-6, with a warm-up ratio of 0.03, while freezing the vision tower during training. Following the default setting for instruction models, the format reward coefficient is set to 0.5. We employ the online filtering strategy with lower and upper bounds of 0.1 and 0.9, respectively. The training process costs 30 hours on 24 H800 GPUs.
A.2 Benchmark Details
@@
ChartQAPro Masry et al. (2025a) introduces a diverse benchmark with various chart types, including infographics and dashboards, and question formats that better reflect real-world challenges. We evaluate models using Chain-of-Thought (CoT) prompting in the original paper and report overall accuracy across five question types.
ChartMuseum Tang et al. (2025a) is a chart question-answering benchmark designed to evaluate complex visual reasoning capabilities with expert-annotated questions from diverse real-world charts. Following the original paper, we evaluate models using the provided CoT prompt and LLM-as-a-Judge evaluation method.
A.3 More Ablation Studies
RL Setting ChartQA CharXiv-RQ
Accuracy Reward
ED 89.88 44.0
ED + SM 90.28 45.2
RL Training Set
QA 89.32 42.1
QA + RQA 90.32 44.6
QA + RQA 90.28 45.2
(a) Ablation study on reward and training set in RL.
SFT Setting ChartQA CharXiv-RQ
SFT Training Set
RQA-SFT&RL + QA 88.40 41.2
RQA-SFT 89.88 44.5
SFT Dataset
TinyChart 84.80 36.1
ChartGemma 86.72 39.1
ChartRQA-SFT 90.20 45.0
(b) Ablation study on training set and different datasets in SFT.
Table 5: Ablation studies within RL and SFT stages. ED and SM are the abbreviations of Edit Distance and Soft Matching. RQA indicates that only samples with chart types of line, bar, and pie in ChartRQA are used for training.
Reward Function To assess different accuracy rewards, we conduct experiments by training Qwen2.5-VL-7B-Instruct for the RL stage only, as shown in Table 5(a). The results demonstrate that employing a soft accuracy reward, which combines edit distance for string-based tasks and soft matching for numerical tasks, yields superior performance across both benchmarks. This finding underscores the importance of adjusting the reward function to the specific type of answers.
Image Diversity & Question Complexity Our ChartRQA dataset is characterized by two key features: diverse chart images and complex questions requiring step-by-step reasoning. To investigate the importance of these factors in RL training, we select samples from ChartRQA-RL that only include line, bar, and pie chart types, which are the same chart types found in the ChartQA dataset, and train the model using RL only. As shown in Table 5(a), without increasing chart type diversity, the complex questions in ChartRQA still substantially enhance the model’s reasoning ability. Furthermore, using the full ChartRQA dataset, which includes a wider variety of chart images, leads to further improvements on CharXiv-RQ.
SFT Data Composition When training Chart-R1, our SFT dataset consists of 228k samples from our ChartRQA-SFT. We then ablate the SFT data composition by adding two sources, the ChartQA dataset and the 30k ChartRQA-RL that overlaps with the RL data, to assess the impact on performance. We train each setting for 2k steps and 1 epoch for SFT and RL, respectively. The results in Table 5(b) show that combining ChartQA and ChartRQA-RL, the final performance decreases evidently. Our analysis indicates that using overlapping data for SFT and RL leads to overfitting, where the model memorizes reasoning paths from the SFT stage, resulting in more rigid thinking processes and a significant loss of output diversity. Also, the direct-answer format of ChartQA discourages the model from developing the ability to break down problems into step-by-step thinking process.
Comparison with Existing Chart Datasets To enable a fair comparison between ChartRQA and existing chart SFT datasets, we replace the SFT dataset with TinyChart and ChartGemma, while keeping all other settings consistent. TinyChart is a comprehensive dataset that integrates multiple open-source datasets and comprises a variety of tasks. To ensure that the model focuses on chart understanding and reasoning, we exclude the Chart-to-text and Chart-to-table generation tasks. For the RL stage, we use a combination of ChartQA and ChartRQA-RL for 1 epoch of training. As shown in Table 5(b), Chart-R1 trained on ChartRQA-SFT achieves the best performance on both benchmarks. The results indicate that the unified thinking and answer format and the effective step-by-step reasoning process in ChartRQA are key factors in enhancing the model’s reasoning ability.
A.4 ChartRQA Analysis
Figure 7: Single-chart samples of 24 chart types from ChartRQA.
Figure 8: Multi-chart samples of 24 chart types from ChartRQA.
We count the quantity and distribution of different chart types across the training and test sets of ChartRQA, as detailed in Table 6. The distribution among the various types to be well-balanced. Furthermore, Figures 7 and  8 provide visualization examples of 24 chart types from the ChartRQA dataset, showcasing both single-chart and multi-chart formats, respectively.
@@ A.4 ChartRQA Analysis
Table 6: The detailed chart types and corresponding quantities in our proposed ChartRQA train and test set. ChartRQA contains 24 chart types, each of which contains approximate samples.
A.5 Prompts
To enhance transparency and reproducibility, we provide the exact prompts used for dataset generation and evaluation. For data generation, we employ Gemini-2.5-Flash to generate both plotting code and QA pairs for data construction. Figure 9 illustrates the prompt used for plotting code generation. We utilize real table data as input, select one chart type from the 24 predefined chart types, and sample a code example corresponding to that chart type to generate the plotting code. Figures 10 and  11 display the prompts used to generate reasoning QA pairs for single-chart and multi-chart formats, respectively. We craft an example for each format to aid LLMs in understanding complex chart reasoning tasks and to generate step-by-step reasoning processes and precise answers that conform to the format. The executable plotting code is provided as auxiliary information to LLMs, making the generated QA pairs more reliable. Figure 12 shows the prompt used for model evaluation. We employ GPT-4o to assess the match between the ground truth and the model’s predictions, where GPT-4o returns a score of 0 or 1 to indicate the correctness of the model’s prediction. Our evaluation focuses solely on the correctness of the final answer, disregarding the reasoning process.
A.6 The Use of AI Assistants
In this work, we used LLMs as writing tools to improve language clarity and readability. These models helped refine the text and enhance the presentation of our ideas. All research concepts, experiments, implementations, and analyses were conducted independently by the authors.
Figure 9: Prompt for code generation.
Figure 10: Prompt for reasoning QA pairs generation for single-chart formats.
Figure 11: Prompt for reasoning QA pairs generation for multi-chart formats.
Figure 12: Prompt for ChartRQA evaluation using GPT-4o.